OpenStudy (anonymous):

If a pair of vertical angles measure 5x-55 and 3x-25, find the angles. I have no idea why I am having such a difficult time with this...Thanks in advance.

OpenStudy (nikato):

WELCOME! And your equation would be 5x-55=3x-25 Because there's a theorem that states that vertical angles are congruent

OpenStudy (anonymous):

@shanef all you need to do is equate both measures (by virtue of Vertical Angle Theorem). 5x-55=3x-25 5x-3x=-25+55 2x=30 x=15 5(15) - 55 = 20 3(15) - 25 = 20

OpenStudy (anonymous):

I feel so stupid. I kept finding "X" correctly, but failed to incorporate "X" into the original equation to find the answer they were actually looking for. Thank you so much. I really appreciate it.

OpenStudy (anonymous):

Pleasure is all mine